# Trace field degrees in the Torelli group

### v1

@@ 6. Explicit pseudo-Anosov maps
The goal of this section is to construct, as explicitly as possible, pseudo-Anosov maps whose stretch factors have prescribed algebraic degrees. In a first step, we build upon our examples in Section 4 to realise also all trace field degrees $d<3g-3$ on $S_{g}$ for $g\geq 2$.
6.1. Multicurve intersection degrees $d<3g-3$
We need a new building block for our surfaces, see Figure 10.
@@ 6.1. Multicurve intersection degrees $d<3g-3$
We denote the red multicurve by $\alpha$ and the blue multicurve by $\beta$. The multicurve $\alpha$ has $k+1$ separating components: one for each of the handles that separates the handle, and one that separates all the handles. We denote the component of $\alpha$ that separates all the handles of the surface in Figure 10 by $\alpha_{1}$, and we denote the other separating components of $\alpha$ by $\alpha_{2},\alpha_{4},\dots,\alpha_{2g-2}$ from left to right. Finally, the remaining nonseparating components of $\alpha$ are $\alpha_{3},\alpha_{5},\dots,\alpha_{2g-1}$ from left to right.
@@
Let $p_{y_{i}}(t)=t^{2}-(4+y_{i})t+4(y_{i}-1)$ be the characteristic polynomial of $B_{y_{i}}$. We know from Section 4 that $p_{y_{i}}$ is irreducible if $y\geq 12$. So, choosing all $y_{i}\geq 12$ pairwise distinct, Remark 11 guarantees that the polynomial $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t][y]$ is irreducible. By Hilbert’s irreducibility theorem, there are infinitely many specifications of $y$ such that $y^{2}-k>0$ and such that $\chi_{XX^{\top}}(t)\in\mathbb{Z}[t]$ is irreducible and of degree $12k+$ 1.
Case 1: $32g\leq d<3g-$ 3. Assume we want to realise the multicurve intersection degree $3g-3-f$ for $30<f\leq g-$ 3. Let $k=f+2\leq g-1$. Start the inductive procedure as in Section 4 with the surface from Figure 10 as a starting point, adding $g-1-k$ more handles. The exact same argument yields a surface of genus $g-1$ with two boundary components, and a characteristic polynomial $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ that is irreducible and of degree $12k+1+3(g-1-k)=3g-3-k+$ 1. Closing up the surface exactly as in Section 4 yields $3g-3-k+2=3g-3-f$ as a multicurve intersection degree on the closed orientable surface of genus $g$.
Case 2: $g<d<2g$. Assume we want to realise the multicurve intersection degree $2g-f$ for $10<f\leq g-$ 1. Take the surface depicted in Figure 10 for $k=g-1$. Now, remove $f$ of the separating curve $\alpha_{2},\dots,\alpha_{2g-2}$. This slightly modifies the matrix $XX^{\top}$$f$ of the 2-by-2 blocks on the diagonal are now 1-by-1 blocks, with the single coefficient $y_{i}$. Nevertheless, since all the $y_{i}$ are chosen pairwise distinct, Remark 11 guarantees that the polynomial $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t][y]$ is irreducible. We note that for the coefficients $y_{i}$ in the 1-by-1 blocks, any positive integer can be chosen. By Hilbert’s irreducibility theorem, there are infinitely many specifications of $y$ such that $\chi_{XX^{\top}}(t)\in\mathbb{Z}[t]$ is irreducible and of degree $2g-1-f$. Closing up the surface as in Section 4 yields the multicurve intersection degree $2g-f$ on the closed orientable surface of genus $g$.
@@ 6.2. Even degree stretch factors
Let $\alpha$ and $\beta$ be an example of a pair of multicurves described in Section 6.1, realising a multicurve intersection degree $31\leq d\leq 3g-$ 3. Then the mapping class $T_{\alpha}\circ T_{\beta}$ is pseudo-Anosov with stretch factor $\lambda$ of degree $2d$.
@@
Proof of 13.
According to 12, all there is to show is
@@
Case 1: $32g\leq d\leq 3g-$ 3. We consider the submatrix $\Omega^{\prime}$ of $\Omega$ that is obtained by deleting all the rows and columns corresponding to components of the multicurve $\alpha$ that have been added during the inductive step or closing up of the surface. Furthermore, if $d<3g-3$, we also remove the component of $\alpha$ encircling multiple handles of the starting surface, that is, the surface depicted in Figure 10.
A base change by a permutation matrix brings $\Omega^{\prime}+2I$ into block diagonal form with $g-1$ blocks corresponding to genus one surface pieces as depicted in Figure 1, and a block of the form $2I$. For a block of the former type, and for $y>4$, we directly calculate that the nullity is zero and the signature equals the dimension of the block minus two. Already, this implies that certainly the signature of $\Omega+2I$ is not equal to its dimension, and it only remains to verify the lower bound in Equation 1.
By construction, if the genus equals $g\geq 2$, we have $g-1$ surface pieces as in Figure 1. This in particular implies that $\sigma(\Omega^{\prime})=\dim(\Omega^{\prime})-2g+2$. Furthermore, we have $\dim(\Omega)-\dim(\Omega^{\prime})=d-2g+2$. The latter equality follows from that fact that the number of components of $\alpha$ in our construction is exactly $d$, and there are two components per surface pieces as in Figure 1. We now calculate
@@
Case 2: $g<d<2g$. We consider the submatrix $\Omega^{\prime}$ of $\Omega$ that is obtained by deleting two rows and two columns corresponding to components of the multicurve $\alpha$: the one corresponding to the component encircling multiple handles in Figure 10 and the one obtained from closing the surface. Recall that we have removed $f=2g-d$ separating curves $\alpha_{2},\dots,\alpha_{2g-2}$.
[…]
@@
By construction, if the genus equals $g\geq 2$, we have $g-1$ surface pieces as in Figure 1. Having at least one piece with $y\leq 3$, this implies that $\sigma(\Omega^{\prime})>\dim(\Omega^{\prime})-2g+2$. Furthermore, we have $\dim(\Omega)-\dim(\Omega^{\prime})=2$. We now calculate
@@
6.3. Explicit examples with stretch factor degree $66g-$ 6
We conclude this section by giving explicit computations supporting a conjecture on the irreducibility of the characteristic polynomials constructed in Section 4 for specific values of $y$. In the inductive step of Section 4, one uses a map
@@ 6.3. Explicit examples with stretch factor degree $66g-$ 6
with the condition $y^{2}>\frac{1}{4}(M_{g})_{11}=(y^{(g)})^{2}$.
By computer assistance, one immediately checks the following proposition.
@@
Together with 13, this gives explicit examples of pseudo-Anosov maps realizing the upper bound $66g-$ 6 in 3 for every $1<g\leq 201$. We don’t know whether $\chi_{M_{g}}$ and $\chi_{N_{g}}$ are actually irreducible for every $g>200$ with the parameters $y^{(i)}=i+1$ chosen as in 14.

### v2

@@ 6. Explicit pseudo-Anosov maps
There are two goals of this section, both inspired by a problem posed by Margalit [Mar19, Problem 10.4]. This section is completely independent of the proof of our main results, 1, 2 and 3.
The first goal is to construct, as explicitly as possible, pseudo-Anosov maps whose stretch factors have prescribed algebraic degrees. For suitably chosen examples, we can replace our nonsplitting criterion, 4, by the nonsplitting criterion of [LL24, Theorem 6]. This provides that the mapping class $T_{\alpha}\circ T_{\beta}$ has degree two over the trace field, which is more explicit than what we obtain using 4. However, we are still not able to control the exact number of parallel copies used for some of the components $\beta_{i}$. This construction is carried out in Section 6.1 and Section 6.2.
The second goal is to construct completely explicit pseudo-Anosov mapping classes of the form $T_{\alpha}\circ T_{\beta}$ with maximal stretch factor degree 6g66g-6. We are able to achieve this for $g\leq 201$ in Section 6.3.
6.1. Multicurve intersection degrees $d\leq 3g-3$
Recall that in Section 4, we constructed pairs of multicurves for which the Thurston–Veech construction realises the multicurve intersection degree 3g33g-3, for $g\geq 2$. In a first step, we build upon our examples from Section 4 to realise also all possible smaller multicurve intersection degrees $1\leq d<3g-3$ on $S_{g}$ for $g\geq 2$.
We need a new building block for our surfaces, which is depicted in Figure 11.
@@ 6.1. Multicurve intersection degrees $d\leq 3g-3$
We denote the red multicurve by $\alpha$ and the blue multicurve by $\beta$. The multicurve $\alpha$ has $k+1$ separating components: one for each of the handles that separates the handle, and one that separates all the handles. We denote the component of $\alpha$ that separates all the handles of the surface in Figure 11 by $\alpha_{1}$, and we denote the other separating components of $\alpha$ by $\alpha_{2},\alpha_{4},\dots,\alpha_{2k}$ from left to right. Finally, the remaining nonseparating components of $\alpha$ are $\alpha_{3},\alpha_{5},\dots,\alpha_{2k+1}$ from left to right.
@@
Let $p_{y_{i}}(t)=t^{2}-(4+y_{i})t+4(y_{i}-1)$ be the characteristic polynomial of $B_{y_{i}}$. We know from Section 4 that $p_{y_{i}}$ is irreducible if $y\geq 12$. So, choosing all $y_{i}\geq 12$ pairwise distinct, Remark 12 guarantees that the polynomial $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t][y]$ is irreducible. By Hilbert’s irreducibility theorem, there are infinitely many specifications of $y$ such that $y^{2}-k>0$ and such that $\chi_{XX^{\top}}(t)\in\mathbb{Z}[t]$ is irreducible and of degree $2k+1$.
Case 1: $2g\leq d<3g-3$. Assume we want to realise the multicurve intersection degree 3g3f3g-3-f for $0<f\leq g-3$. Let $k=f+2\leq g-1$. Start the inductive procedure as in Section 4 with the surface from Figure 11 as a starting point, adding g1kg-1-k more handles. The exact same argument yields a surface of genus g1g-1 with two boundary components, and a characteristic polynomial $\chi_{XX^{\top}}\in\mathbb{Z}[t]$ that is irreducible and of degree 2k+1+3(g1k)=3g3k+12k+1+3(g-1-k)=3g-3-k+1. Closing up the surface exactly as in Section 4 yields 3g3k+2=3g3f3g-3-k+2=3g-3-f as a multicurve intersection degree on the closed orientable surface of genus $g$.
Case 2: $g<d<2g$. Assume we want to realise the multicurve intersection degree 2gf2g-f for $0<f\leq g-1$. Take the surface depicted in Figure 11 for k=g1k=g-1. Now, remove $f$ of the separating curve $\alpha_{2},\dots,\alpha_{2g-2}$. This slightly modifies the matrix $XX^{\top}$$f$ of the 2-by-2 blocks on the diagonal are now 1-by-1 blocks, with the single coefficient $y_{i}$. Nevertheless, since all the $y_{i}$ are chosen pairwise distinct, Remark 12 guarantees that the polynomial $\chi_{XX^{\top}}(t,y)\in\mathbb{Z}[t][y]$ is irreducible. We note that for the coefficients $y_{i}$ in the 1-by-1 blocks, any positive integer can be chosen. By Hilbert’s irreducibility theorem, there are infinitely many specifications of $y$ such that $\chi_{XX^{\top}}(t)\in\mathbb{Z}[t]$ is irreducible and of degree 2g1f2g-1-f. Closing up the surface as in Section 4 yields the multicurve intersection degree 2gf2g-f on the closed orientable surface of genus $g$.
@@ 6.2. Even degree stretch factors
Let $\alpha$ and $\beta$ be an example of a pair of multicurves described in Section 6.1, realising a multicurve intersection degree $1\leq d\leq 3g-3$. Then the mapping class $T_{\alpha}\circ T_{\beta}$ is pseudo-Anosov with stretch factor $\lambda$ of degree $2d$.
@@
Proof of 14.
According to 13, all there is to show is
@@
Case 1: $2g\leq d\leq 3g-3$. We consider the submatrix $\Omega^{\prime}$ of $\Omega$ that is obtained by deleting all the rows and columns corresponding to components of the multicurve $\alpha$ that have been added during the inductive step or closing up of the surface. Furthermore, if d<3g3d<3g-3, we also remove the component of $\alpha$ encircling multiple handles of the starting surface, that is, the component $\alpha_{1}$ of the surface depicted in Figure 11.
A base change by a permutation matrix brings $\Omega^{\prime}+2I$ into block diagonal form with g1g-1 blocks corresponding to genus one surface pieces as depicted in Figure 1, and a block of the form $2I$. For a block of the former type, and for $y>4$, we directly calculate that the nullity is zero and the signature equals the dimension of the block minus two. Already, this implies that certainly the signature of $\Omega+2I$ is not equal to its dimension, and it only remains to verify the lower bound in Equation 1.
By construction, if the genus equals $g\geq 2$, we have g1g-1 surface pieces as in Figure 1. This in particular implies that $\sigma(\Omega^{\prime})=\dim(\Omega^{\prime})-2g+2$. Furthermore, we have $\dim(\Omega)-\dim(\Omega^{\prime})=d-2g+2$. The latter equality follows from that fact that the number of components of $\alpha$ in our construction is exactly $d$, and there are two components per surface pieces as in Figure 1. We now calculate
@@
Case 2: $g<d<2g$. We consider the submatrix $\Omega^{\prime}$ of $\Omega$ that is obtained by deleting two rows and two columns corresponding to the following components of the multicurve $\alpha$: the one corresponding to the component $\alpha_{1}$ encircling multiple handles in Figure 11 and the one obtained from closing the surface. Recall that we have removed f=2gdf=2g-d of the separating curves $\alpha_{2},\dots,\alpha_{2g-2}$.
A base change by a permutation matrix brings $\Omega^{\prime}+2I$ into block diagonal form with g1fg-1-f blocks corresponding to surface pieces as in Figure 1ff blocks corresponding to surface pieces as in Figure 1 but with the separating component of $\alpha$ removed, and a block of the form $2I$.
@@
By construction, if the genus equals $g\geq 2$, we have g1g-1 surface pieces as in Figure 1. Having at least one piece with $y\leq 3$, this implies that $\sigma(\Omega^{\prime})>\dim(\Omega^{\prime})-2g+2$. Furthermore, we have $\dim(\Omega)-\dim(\Omega^{\prime})=2$. We now calculate
@@
6.3. Explicit examples with stretch factor degree 6g66g-6
We conclude this section by giving explicit computations, for $1<g\leq 201$, supporting a conjecture on the irreducibility of the characteristic polynomials constructed in Section 4 for specific values of $y$. In the inductive step of Section 4, one uses a map
@@ 6.3. Explicit examples with stretch factor degree 6g66g-6
with the condition $y^{2}>\frac{1}{4}(M_{g})_{11}=(y^{(g-1)})^{2}$.
By computer assistance [S+YY], one immediately checks the following proposition.
@@
Together with 14, this gives explicit examples of pseudo-Anosov maps realising the upper bound 6g66g-6 in 3 for every $1<g\leq 201$. We do not know whether $\chi_{M_{g}}$ and $\chi_{N_{g}}$ are actually irreducible for every $g>200$ with the parameters $y^{(i)}=i+1$ chosen as in 15.